# Chiral excitation flows of multinode network based on synthetic gauge fields

### v1

@@ III.1 Perfect chiral flow assisted with auxiliary node in a four-node network
In short, we conclude that the sub-triangle structure’s TR symmetry determines the dynamical chirality of the bosonic multinode network. When the additional auxiliary node is introduced to couple to network nodes and the synthetic flux of sub-triangles $\theta_{\triangle}$ is chosen to be the phase referred to in the three-node case, i.e., $(2\mathbb{Z}+1)\pi/2$, the perfect chiral flow can emerge with a specific value of $\beta_{c}$. In particular, the $\pi/2$ hopping phase and the auxiliary node are the two essential elements for achieving the perfect chiral flow in the four-node boson-aSGF model. But as we directly extend the boson-aSGF model to $n$-node networks, due to the anharmonicity of energy spectrum, the chiral flow collapses regardless of the strength of $\beta_{c}$, as detailed in Appendix B, and explicit expressions for six-node aSGF model can be found in Appendix C. Nevertheless, the study of four-node network demonstrates the principle of conducting chiral flows from the quantum interference of different states. In the following subsection, we will show that, by introducing the next-nearest-neighbor (NNN) hopping in five- and six-node networks, perfect chiral flows can revive with the boson-aSGF model.
@@ III.2 Perfect $n$-node chiral flow beyond NN interactions
leading to a symmetric energy spectrum of the system with paired eigenstates. Owing to the disparity in zero-energy eigenstates, we separately analyze the scenarios for even and odd values of $n$. With identical couplings between the auxiliary node and $n$ network nodes, the Hamiltonian can be written as
@@
where $|\varphi_{m}\rangle~{}(-n/2<m\leq n/2)$ are the eigenstates of $\mathcal{H}_{\mathrm{SGF}}^{n}$, as discussed in Appendix B, and $|c\rangle=c^{\dagger}|G\rangle$ is the excited state of the auxiliary node. Here, the energy spectrum $E_{m}=-2J_{0}\sin k_{m}$ exhibits the dispersion relation of one-dimensional tight-binding Hamiltonian, where $k_{m}=2\pi m/n$ are wave numbers in the first Brillouin zone. We refer to these eigenstates $|\varphi_{m}\rangle$ as chiral modes in this system because they have chiral phase shifts $2\pi m/n$ between nearest sites. For an odd number of $n$, the zero-energy state $|\varphi_{0}\rangle$ hybridize with the product state $|c\rangle$, and they eventually form the new paired chiral modes $|\varphi_{\beta_{c},\pm}\rangle=(|\varphi_{0}\rangle\pm|c\rangle)/\sqrt{2}$, which are independent of the value of $\beta_{c}$, while the eigenenergies ±$\pm E_{\beta_{c}}=\pm\sqrt{n}\beta_{c}J_{0}$ are tuned by $\beta_{c}$. If $n$ is an even number, the chiral symmetry of $H_{\mathrm{SGF}}^{n}$ ensures the existence of two degenerate zero-energy states, corresponding to $|\varphi_{0}\rangle$ and $|\varphi_{n/2}\rangle$. Similarly, the auxiliary node results in the hybridization between the two states $|\varphi_{0}\rangle$ and $|c\rangle$, and thus the degeneracy is lifted. Instead, $|\varphi_{n/2}\rangle$ remains as the unperturbed zero-energy state. We emphasize that the hybridized states together with those unperturbed states $|\varphi_{m}\rangle$ form a complete set of required chiral modes, which makes it possible to realize a perfect chiral flow in this $n$-node network.
Figure 3: Dynamics of the modified (a) five-node and (b) six-node aSGF models. The NNN hopping is introduced in the modified networks and ensures perfect chiral flows.
We can now deduce the Hamiltonian from the prerequisites to create a perfect chiral flow in the $n$-node network. Here, we illustrate the methodology for constructing chiral flow Hamiltonians from the boson-aSGF model by considering the examples of five- and six-node networks, corresponding to odd and even values of $n$, respectively, where the initial condition is still denoted by an excitation localized on the 1-st node. We first consider the five-node case, a paradigm for the odd $n$. Since the eigenstates are independent of the coupling strength $\beta_{c}$, the excitation amplitude for $j$-th node is given by
@@
Therefore, conducting a perfect chiral flow requires the complete transmission condition: $\max[|C_{e}^{j}(t)|^{2}]=1$ for all sites and thus leads to the conditions $E_{2}=2E_{1}$ and
@@
which also reduces to $E_{\beta_{c}}=5E_{1}$. According to the eigenstates and these constraints on eigenenergies, we yield the chiral flow Hamiltonian from this five-node boson-aSGF model, with the NNN hopping:
@@
where $\alpha=\sqrt{(3-\sqrt{5})/2},~{}\beta_{c}=5\sqrt{2/(5+\sqrt{5})}$. As shown in Fig. 3(a), we plot the dynamical excitation flow with Hamiltonian $H_{5}$. Here, for a clockwise chiral flow, the NN hopping phase is $-\pi/2$ instead of $\pi/2$. Both the NN and NNN hoppings exhibit a phase of $(\mathbb{Z}+1\pi)/2$ to preserve the chiral symmetry.
@@
where the factor $(-1)^{j-1}$ is interpreted as $\cos(j-1)\pi$. Hence, the eigenstates $|\varphi_{\pm m}\rangle$ correspond to energies $\mp E_{3-m}$. From Eq. (13) we can derive the energy-level relationships as follows: $E_{1}=2E_{2},~{}E_{\beta_{c}}=3E_{2}$. With the eigenstates and energy levels, we obtain the six-node chiral flow Hamiltonian
@@
In Fig. 3(b), we show the perfect chiral flow in the six-node network assisted by the auxiliary node. In this sense, we have found a universal scheme for implementing the chiral flow within both odd and even categories of $n$, and when we go to the $n$-node boson-aSGF network, there are more interactions involved, with Hamiltonian given by the form
$H_{n}/J_{0}=\sum_{j=1}^{n}\sum_{k=1}^{\lceil n/2\rceil-1}i\alpha_{k}a_{j}^{\dagger}a_{j+k}+\beta_{c}a_{j}^{\dagger}c+\mathrm{H.c.},$ (15)
@@
However, as the number of nodes $n$ increases, the complexity of the aSGF network for conducting a perfect chiral flow also escalates. Specifically, the number of required couplings exhibits a scaling of $\sim n^{2}$, which results in the difficulty of constructing a chiral flow in practical systems. Thus, we will turn to find an alternative way to realize chiral transmission between remote nodes in a large network in Sec. IV.
III.3 Spin-aSGF model and multi-excitation
Figure 4: (a) Schematics of the three-node spin-SGF model. TR symmetry predicts opposite chiral flows for the spin-up state and spin-down state. (b) The time evolution of Bell states when the hopping phase is $\pi/2$. The Bell states $|\Psi_{jk}^{+}\rangle$ and $|\Phi_{jk}^{+}\rangle$ corresponding to different excitation subspace have opposite chiral dynamics, where the concurrence of two atoms $\mathscr{C}_{jk}$ transfers with the evolution of population $P(|\Psi_{jk}^{+}\rangle)$ or $P(|\Phi_{jk}^{+}\rangle)$.
In the two-level atomic and spin systems, we can also develop the spin-SGF model by synthesizing the complex hoppings between atoms [52, 53] or the spin interaction [23]. However, the spin-SGF model exhibits very different symmetries compared to the boson-SGF model, resulting in a distinctive behavior of its chiral dynamics. By replacing the boson annihilation and creation operators $a_{j},~{}a_{j}^{\dagger}$ with ladder operators $\sigma_{j}^{-},\sigma_{j}^{+}$, we obtain the spin-SGF Hamiltonian
@@ III.3 Spin-aSGF model and multi-excitation
This tight-binding Hamiltonian can describe both the excitation hopping between two-level atoms and the spin-exchange interaction. Different from the description in spinless lattice models [10], the TR operator for the spin system is defined as $\hat{T}=i\sigma_{y}\mathcal{K}$, where $\mathcal{K}$ denotes the complex conjugation. The operator $\hat{T}$ flips spin vector as
@@
Furthermore, based on the unique dynamics in the spin-SGF network, we can establish opposite chiral transmission channels for different entanglement states. In Fig. 4(b), we illustrate the opposite chiral dynamics of two types of Bell states $|\Psi^{+}_{jk}\rangle=(\left|\uparrow_{j}\downarrow_{k}\right\rangle+\left|\downarrow_{j}\uparrow_{k}\right\rangle)/\sqrt{2}$ and $|\Phi^{+}_{jk}\rangle=(\left|\downarrow_{j}\downarrow_{k}\right\rangle+\left|\uparrow_{j}\uparrow_{k}\right\rangle)/\sqrt{2}$, which correspond to the spin-up and spin-down chiralities. Here, we set the initial states as $|\Psi_{12}^{+}\rangle\otimes\left|\downarrow_{3}\right\rangle$ and $|\Phi_{12}^{+}\rangle\otimes\left|\downarrow_{3}\right\rangle$ in the left and right panels of Fig. 4(b), respectively. Both the populations of the Bell states and the concurrence $\mathscr{C}_{jk}$ can characterize the chirality of entanglement in this spin network. Because the dynamics of the two Bell states is constrained in the up and down subspaces, the preserved TR symmetry leads to opposite flow directions for them and causes distinctive entanglement flows. However, in the boson network with broken TR symmetry, the chirality remains consistent for different excitation subspaces. Hence, the transmission of entanglement is not sensitive to its initial state. Notably, the three-node case of the spin-SGF Hamiltonian, with the $\pi/2$ hopping phase, is also equivalent to the chiral antisymmetric spin exchange interaction (ASI) Hamiltonian [59, 60], which breaks parity symmetry but preserves TR symmetry [23].
@@
Table 1: A review of different models supporting the three-node chiral flow.
@@
chiral symmetry
Similarly, the $n$-node chiral flow can also be created in the spin or atomic system by employing the aSGF model with the replacing of operators $\sigma^{-(+)}\rightarrow a^{(\dagger)}$ in Eq. (15). As shown in Fig. 5(a), we observe chiral dynamics of the spin-up state in the four-node spin-aSGF network. If one then flips all the spins, an opposite chiral flow will arise between the network nodes, predicted by TR symmetry, as shown in Fig. 5(b). However, in contrast to the boson system, the spin-aSGF model does not support multi-excitation chiral flows and only permits the chiral dynamics for a single up or down state. In boson-aSGF networks, the unidirectional chiral dynamics exists in each excitation subspace. Therefore, it can be utilized to transfer or prepare more complicated entangled states [23], such as the GHZ state and the NOON state [55].
@@ III.4 Discussion of the criteria
From the study of constructing the Hamiltonian (15) for achieving the $n$-node chiral flow, we conclude that there are two essential ingredients for inducing perfect chiral flows: (i) the symmetric and equally spaced energy spectrum in a specific excitation subspace and (ii) a complete set of the paired chiral modes. The symmetric spectrum is usually obtained from the chiral symmetry of Hamiltonians in previous models, and it guarantees that each set of paired chiral modes with opposite energies contributes a propagating wave of the excitation in the network. Then, the equally spaced energy levels facilitate the interference of all pairs of chiral modes on each site, leading to a perfect site-by-site chiral flow. In particular, the chiral symmetry is not a necessity, as we can only generate the chiral flow in a specific subspace. For example, when we add on-site potential $U$ to the three-node boson chiral flow Hamiltonian such that $H^{\prime}=\sum_{j=1}^{3}U\hat{n}_{j}^{2}+(iJ_{0}a_{j}^{\dagger}a_{j+1}+\mathrm{H.c.})$, the chiral symmetry is broken. However, there is still a symmetric spectrum in the single-excitation subspace, allowing for the existence of the chiral flow. As the potential $U\gg J_{0}$, the energy spectrum in the double-excitation subspace also becomes nearly symmetric. In this sense, it forms a hard-core boson system and mimics the three-node spin-SGF network, resulting in opposite chiral flows in these two subspaces [22].
@@
On the other hand, additional symmetries can also impose constraints on the chiral flow Hamiltonian and bring novel phenomena, e.g., the TR symmetry of the three-node spin-SGF model promotes different chirality like the quantum spin Hall effect. As shown in Table 1, we summarize different three-node models that support perfect chiral flows, and all of them manifest the chiral symmetry mentioned above. These models may break other symmetries within their systems, such as parity symmetry [23] (described by $\hat{P}$ operator) and TR symmetry, while the spin chirality interaction (SCI) Hamiltonian [61, 62, 24] breaks both of them. In particular, all the Hamiltonians of the listed models have the $U(1)$ symmetry that preserves the excitation number. Therefore, we can block-diagonalize the Hamiltonians as $H=\mathcal{H}^{(0)}\otimes\mathcal{H}^{(1)}\otimes\cdots\otimes\mathcal{H}^{(n)}$, where $H^{(j)}$ is the block Hamiltonian in the $j$-excitation subspace (for spin system the excitation is replaced by spin-up state). The symmetric energy spectrum of each block Hamiltonian makes the chiral flow present in each subspace. In our aSGF model, the chiral symmetry is still preserved after introducing an auxiliary node if the hopping (including long-range hoppings in the $n$-node case) phase between network nodes is $\theta_{j,j+k}=(2\mathbb{Z}+1)\pi/2$. The auxiliary node also generates the last required pair of chiral modes, which facilitates the completeness of chiral modes and thus re-induces perfect chiral flows in the $n$-node aSGF networks.

### v2

@@ III.1 Perfect chiral flow assisted by an auxiliary node in a four-node network
In short, we conclude that the subtriangle structure’s TR symmetry determines the dynamical chirality of the bosonic multinode network. When the additional auxiliary node is introduced to couple to network nodes, and the synthetic flux of subtriangles $\theta_{\triangle}$ is chosen to be the phase referred to in the three-node case, i.e., $(p+1/2)\pi$ with $p\in\mathbb{Z}$, the perfect chiral flow can emerge with a specific value of $\beta_{c}$. In particular, the $\pi/2$ hopping phase and the auxiliary node are the two essential elements for achieving the perfect chiral flow in the four-node boson-aSGF model; however, when we directly extend the boson-aSGF model to $n$-node networks, due to the anharmonicity of energy spectrum, the chiral flow collapses regardless of the strength of $\beta_{c}$ (as detailed in Appendix B), and explicit expressions for a six-node aSGF model can be found in Appendix C. Nevertheless, the study of the four-node network demonstrates the principle of conducting chiral flows from the quantum interference of different states. In the next subsection, we will show that by introducing the next-nearest-neighbor (NNN) hopping in five- and six-node networks, perfect chiral flows can be revived with the boson-aSGF model.
@@ III.2 Perfect $n$-node chiral flow beyond NN interactions
leading to a symmetric energy spectrum for the system with paired eigenstates. Due to the disparity in zero-energy eigenstates, we separately analyze the scenarios for even and odd values of $n$. With identical couplings between the auxiliary node and $n$ network nodes, the Hamiltonian can be written as
@@
where $|\varphi_{m}\rangle~{}(-n/2<m\leq n/2)$ are the eigenstates of $\mathcal{H}_{\mathrm{SGF}}^{n}$, as discussed in Appendix B, and $|c\rangle=c^{\dagger}|G\rangle$ is the excited state of the auxiliary node. Here, the energy spectrum $E_{m}=-2J_{0}\sin k_{m}$ exhibits the dispersion relation of the one-dimensional tight-binding Hamiltonian, where $k_{m}=2\pi m/n$ are wave numbers in the first Brillouin zone. We refer to these eigenstates $|\varphi_{m}\rangle$ as chiral modes in this system because they have chiral phase shifts $2\pi m/n$ between nearest sites. For an odd value of $n$, the zero-energy state $|\varphi_{0}\rangle$ hybridizes with the product state $|c\rangle$, and they eventually form the new paired chiral modes $|\varphi_{\beta_{c},\pm}\rangle=(|\varphi_{0}\rangle\pm|c\rangle)/\sqrt{2}$, which are independent of the value of $\beta_{c}$, while the eigenenergies ±$\pm E_{\beta_{c}}=\pm\sqrt{n}\beta_{c}J_{0}$ are tuned by $\beta_{c}$. If $n$ is an even number, the chiral symmetry of $H_{\mathrm{SGF}}^{n}$ ensures the existence of two degenerate zero-energy states, corresponding to $|\varphi_{0}\rangle$ and $|\varphi_{n/2}\rangle$. Similarly, the auxiliary node results in hybridization between the two states $|\varphi_{0}\rangle$ and $|c\rangle$, and the degeneracy is thus lifted. Instead, $|\varphi_{n/2}\rangle$ remains as the unperturbed zero-energy state. We emphasize that the hybridized states, together with those unperturbed states $|\varphi_{m}\rangle$, form a complete set of required chiral modes, which makes it possible to realize a perfect chiral flow in $n$-node networks.
Figure 3: Dynamics of the modified (a) five-node and (b) six-node aSGF models. NNN hopping is introduced in the modified networks to ensure perfect chiral flows.
We now deduce the Hamiltonian from the prerequisites to create a perfect chiral flow in the $n$-node network. Here, we illustrate the methodology for constructing chiral-flow Hamiltonians from the boson-aSGF model by considering the examples of five- and six-node networks, corresponding to odd and even values of $n$, respectively, where the initial condition is still denoted by an excitation localized on the first node. The Hamiltonian realizing chiral flow in the $n$-node ($n>4$) network is labeled by $H_{\mathrm{ch}}^{n}$, as it further involves NNN hoppings in comparison to the simple aSGF Hamiltonian $H_{\mathrm{aSGF}}^{n}$. We first consider the five-node case, a paradigm for odd $n$. Since the eigenstates are independent of the coupling strength $\beta_{c}$, the excitation amplitude for $j$th node is given by
@@
Therefore, conducting a perfect chiral flow requires the complete transmission condition: $\max[|C_{e}^{j}(t)|^{2}]=1$ for all sites, leading to the conditions $E_{2}=2E_{1}$ and
@@
which also reduces to $E_{\beta_{c}}=5E_{1}$. According to the eigenstates and these constraints on eigenenergies, we yield the following chiral-flow Hamiltonian from this five-node boson-aSGF model, with NNN hopping:
@@
where $\alpha=\sqrt{(3-\sqrt{5})/2},~{}\beta_{c}=5\sqrt{2/(5+\sqrt{5})}$. In Fig. 3(a), we plot the dynamical excitation flow with Hamiltonian $H_{\mathrm{ch}}^{5}$. Here, for a clockwise chiral flow, the NN hopping phase is $-\pi/2$ instead of $\pi/2$. Both the NN and NNN hoppings exhibit a phase of $(p+1/2)\pi$ with $p\in\mathbb{Z}$ to preserve the chiral symmetry.
@@
where the factor $(-1)^{j-1}$ is interpreted as $\cos(j-1)\pi$. Hence, the eigenstates $|\varphi_{\pm m}\rangle$ correspond to energies $\mp E_{3-m}$. From Eq. (13), we can derive the energy-level relationships as follows: $E_{1}=2E_{2}$ and $E_{\beta_{c}}=3E_{2}$. With the eigenstates and energy levels, we obtain the six-node chiral-flow Hamiltonian
@@
In Fig. 3(b), we show the perfect chiral flow in the six-node network assisted by the auxiliary node. The six-node aSGF network with perfect chiral flow exhibits a total flux of zero, while breaking TR symmetry in its subtriangle structure. This is analogous to a plaquette in the Haldane model [68], a model that demonstrates chiral edge currents without magnetic flux.
By investigating networks with odd and even numbers of nodes, we have found a universal scheme for implementing the chiral flow within both odd and even categories of $n$. When we go to the $n$-node boson-aSGF network, there are more interactions involved, with the Hamiltonian having the form
$H_{\mathrm{ch}}^{n}/J_{0}=\sum_{j=1}^{n}\sum_{k=1}^{\lceil n/2\rceil-1}i\alpha_{k}a_{j}^{\dagger}a_{j+k}+\beta_{c}a_{j}^{\dagger}c+\mathrm{H.c.},$ (15)
@@
III.3 Spin-aSGF model and multiexcitation
Figure 4: (a) Schematic of the three-node spin-SGF model. TR symmetry predicts opposite chiral flows for the spin-up and spin-down states. (b) The time evolution of Bell states when the hopping phase is $\pi/2$. The Bell states $|\Psi_{jk}^{+}\rangle$ and $|\Phi_{jk}^{+}\rangle$ corresponding to different excitation subspaces have opposite chiral dynamics, where the concurrence of two spins (or atoms) $\mathscr{C}_{jk}$ transfers with the evolution of population $P(|\Psi_{jk}^{+}\rangle)$ or $P(|\Phi_{jk}^{+}\rangle)$.
In the two-level atomic and spin systems, we can also develop the spin-SGF model by synthesizing the complex hoppings between atoms [59, 60] or the spin interaction [30]; however, the spin-SGF model exhibits very different symmetries compared to the boson-SGF model, resulting in distinctive chiral dynamics. By replacing the boson annihilation and creation operators $a_{j}$ and $a_{j}^{\dagger}$ with ladder operators $\sigma_{j}^{-}$ and $\sigma_{j}^{+}$, we obtain the spin-SGF Hamiltonian
@@ III.3 Spin-aSGF model and multiexcitation
This tight-binding Hamiltonian can describe both the excitation hopping between two-level atoms and the spin-exchange interaction. In contrast to the description in spinless lattice models [10], the TR operator for the spin system is defined as $\hat{T}=i\sigma_{y}\mathcal{K}$, where $\mathcal{K}$ denotes the complex conjugation. The operator $\hat{T}$ flips the spin vector as
@@
Furthermore, based on the unique dynamics in the spin-SGF network, we can establish opposite chiral transmission channels for different entanglement states. In Fig. 4(b), we illustrate the opposite chiral dynamics of two types of Bell states $|\Psi^{+}_{jk}\rangle=(\left|\uparrow_{j}\downarrow_{k}\right\rangle+\left|\downarrow_{j}\uparrow_{k}\right\rangle)/\sqrt{2}$ and $|\Phi^{+}_{jk}\rangle=(\left|\downarrow_{j}\downarrow_{k}\right\rangle+\left|\uparrow_{j}\uparrow_{k}\right\rangle)/\sqrt{2}$, which correspond to the spin-up and spin-down chiralities. Here, we set the initial states as $|\Psi_{12}^{+}\rangle\otimes\left|\downarrow_{3}\right\rangle$ and $|\Phi_{12}^{+}\rangle\otimes\left|\downarrow_{3}\right\rangle$ in the left- and right-hand panels of Fig. 4(b), respectively. Both the populations of the Bell states and the concurrence $\mathscr{C}_{jk}$ can characterize the chirality of entanglement in this spin network. Because the dynamics of the two Bell states are constrained in the up and down subspaces, respectively, the preserved TR symmetry leads to opposite flow directions for them and causes distinctive entanglement flows; however, in the bosonic network with broken TR symmetry, the chirality remains consistent for different excitation subspaces. Hence, the transmission of entanglement is not sensitive to its initial state. Notably, the three-node case of the spin-SGF Hamiltonian, with the $\pi/2$ hopping phase, is also equivalent to the chiral antisymmetric spin exchange interaction (ASI) Hamiltonian [69, 70], which breaks parity symmetry but preserves TR symmetry [30].
@@
Table 1: Review of different models supporting the three-node chiral flow.
@@
Symmetric spectrum
Similarly, the $n$-node chiral flow can also be created in a spin or atomic system by employing the aSGF model by replacing the operators $a^{(\dagger)}\rightarrow\sigma^{-(+)}$ in Eq. (15). As shown in Fig. 5(a), we observe chiral dynamics of the spin-up state in the four-node spin-aSGF network. If one then flips all the spins, an opposite chiral flow will arise between the network nodes, predicted by TR symmetry, as shown in Fig. 5(b); however, in contrast to the bosonic system, the spin-aSGF model does not support multiexcitation chiral flows and only permits the chiral dynamics for a single up or down state. In boson-aSGF networks with multiexcitation, the unidirectional chiral dynamics exists in each excitation subspace. Since the Hamiltonian does not contain photon-photon interactions, these excitations do not interact with each other and will display the same chiral motion irrespective of the excitation number. Therefore, it can be used to transfer or prepare more complicated entangled states [30], such as the Greenberger-Horne-Zeilinger state and the NOON state [62].
@@ III.4 Discussion of the criteria
From the study of constructing the Hamiltonian (15) for achieving the $n$-node chiral flow, we conclude that there are two essential ingredients for inducing perfect chiral flows: (i) a symmetric and equally spaced energy spectrum in a specific excitation subspace and (ii) a complete set of the paired chiral modes. The symmetric spectrum is usually obtained from the chiral symmetry of Hamiltonians in previous models, and it guarantees that each set of paired chiral modes with opposite energies contributes a propagating wave of the excitation in the network. The equally spaced energy levels then facilitate the interference of all pairs of chiral modes on each site, leading to a perfect site-by-site chiral flow. In particular, chiral symmetry of the whole Hamiltonian is not a prerequisite because we can build a symmetric spectrum and perform the perfect chiral flow exclusively within a specific subspace or several subspaces. For example, when we add the on-site potential $U$ (photon-photon interactions) to the three-node bosonic chiral-flow Hamiltonian such that $H^{\prime}=\sum_{j=1}^{3}U\hat{n}_{j}^{2}+(iJ_{0}a_{j}^{\dagger}a_{j+1}+\mathrm{H.c.})$, the chiral symmetry is broken; however, there is still a symmetric spectrum in the single-excitation subspace, allowing for the existence of the chiral flow. As the potential $U\gg J_{0}$, the energy spectrum in the double-excitation subspace also becomes nearly symmetric. In this sense, it forms a hard-core bosonic system and mimics the three-node spin-SGF network, resulting in opposite chiral flows in these two subspaces [29].
@@
Additional symmetries can also impose constraints on the chiral-flow Hamiltonian and produce previously unseen phenomena, e.g., the TR symmetry of the three-node spin-SGF model promotes different chirality, like the quantum spin Hall effect. In Table 1, we summarize different three-node models that support perfect chiral flows, and all of them manifest the chiral symmetry mentioned above. These models may break other symmetries within their systems, such as parity symmetry [30] (described by $\hat{P}$ operator) and TR symmetry, while the spin-chirality interaction (SCI) Hamiltonian [71, 72, 31] breaks both of them. In particular, all the Hamiltonians of the listed models have $U(1)$ symmetry, which preserves the excitation number. Therefore, we can block-diagonalize the Hamiltonians as $H=\mathcal{H}^{(0)}\otimes\mathcal{H}^{(1)}\otimes\cdots\otimes\mathcal{H}^{(n)}$, where $\mathcal{H}^{(j)}$ is the block Hamiltonian in the $j$-excitation subspace (for a spin system the excitation is replaced by the spin-up state). For each model listed in Table 1, the symmetric energy spectrum of the block Hamiltonian $\mathcal{H}^{(j)}$ makes it possible for chiral flows to be present in the $j$th excitation subspace. In our aSGF model, the chiral symmetry is still preserved after introducing an auxiliary node if the hopping (including long-range hoppings in the $n$-node case) phase between network nodes is $\theta_{j,j+k}=(p+1/2)\pi$ with $p\in\mathbb{Z}$. The auxiliary node also generates the last required pair of chiral modes, which facilitates the completeness of chiral modes and thus restores perfect chiral flows within $n$-node aSGF networks.